# Scaling Laws of Synthetic Data for Language Models

### v2

@@ 5.2 Main Results
SynthLLM consistently outperforms other synthetic datasets across all benchmarks, demonstrating the effectiveness of our framework. OpenMathInstruct-2, a high-quality math reasoning dataset generated using the Llama3.1-405B model based on MATH and GSM8K, comprises 14M samples. Despite its significantly larger size, the model trained on smaller dataset (SynthLLM-8B (3.2M)) achieves comparable performance on its in-distribution test sets (MATH and GSM8K) while substantially surpassing OpenMathInstruct-2 on Out-of-Distribution (OOD) test sets. This highlights the superior generalization capabilities of SynthLLM. Compared to the corresponding Llama Instruct models across different sizes, our models consistently achieve superior performance. Furthermore, our models (SynthLLM-8B) match or even surpass much larger models such as NuminaMath-CoT-72B and Llama-3.1-70B-Instruct on benchmarks, further reinforcing the effectiveness of our approach.
@@
More importantly, for each reference document, we currently generate only five questions per document for Level-1 and Level-2, and three questions for Level-3, which is far from reaching the full potential of our method’s document utilization (as shown in Figure 6). We can further enhance performance on benchmarks simply by increasing the number of synthesized questions per document.
@@ 5.3 Ablation Studies about Our SynthLLM Framework
In this section, we present detailed ablation studies about our Level-2 and Level-3 methods, demonstrating their advantages in scalability and diversity.
@@
Table 3: Performance of 1B and 3B models on various math benchmarks. The number in ()()( ) represents the sample size. The best result is in bolded. All metrics are measured in percentage (%).
Comparison of Level-2 and Level-3 with directly extraction-based synthesis method.
We compare the directly extraction-based synthesis method (Level-1) with our Level-2 and Level-3 methods across models of varying sizes. The results are shown in Table 3. We observe that, on the 3B model, Level-2 and Level-3 significantly outperform Level-1 and instruct version across benchmarks. On the 1B model, Level-3 surpasses Level-1 by a large margin on most benchmarks, with the exception of College and Gaokao. Although each document is sampled with 5 questions, the sample size for Level-1 (2.3M) is significantly lower than that of Level-2 (2.6M).
This is because Level-1 is based on direct extraction, which leads to a substantial amount of duplicate questions being generated. After removing duplicates, the sample size decreases considerably. This also highlights the scalability limitations of direct extraction-based synthesis, as it is inherently constrained by the availability of high-quality reference documents containing questions.
Diversity of Generated Questions. Compared to direct extraction-based synthesis, SynthLLM generates more diverse questions by decomposing and recombining knowledge concepts. To assess this, we compare Level-1 and Level-2 by measuring the similarity among generated questions within the same document. Specifically, we randomly select 2,000 documents and generate 50 questions per document for each method. We adopt all-MiniLM-L6-v2 [49] to obtain question embeddings and compute cosine similarity between each pair of questions within the same document. The histogram of question similarity is shown in Figure 5. We could clearly observe that questions generated by Level-2 exhibit lower similarity, indicating greater diversity among questions derived from the same document.
Figure 6: (a): The performances of Level-2, Level-3, and other augmentation methods on MATH benchmark. (b): the average performances across various benchmarks. The x axis denotes the sample number. The y axis represents the accuracy.
Comparison with other data augmentation methods on limited reference documents.
We compare our Level-2 and Level-3 methods with mainstream augmentation techniques, including rephrase [53] and persona augmentation [12]. We restrict the number of reference documents to a very small set and evaluate different methods for scaling high-quality question generation from a constrained document pool. The effectiveness of each method will be measured by its ability to consistently improve model performance on downstream tasks as the dataset expands. To achieve this goal, We randomly select 2,000 documents and generate 25, 50, 75, …, 150 questions per document—a significantly large number—using both Level-2 and Level-3 methods. This results in datasets of 50K, 100K, 150K, …, 300K questions for each method. For rephrase and persona augmentation methods, we first use Level-1 to extract existing questions from documents, yielding approximately 6,500 questions as seed dataset. We then expand this dataset to 300K questions using these two augmentation techniques. We train these datasets on Llama-3.1-8B model, with results presented in Figure 6. Figure 6 (a) illustrates MATH accuracy of different methods, while Figure 6 (b) shows average accuracy across benchmarks. Level-2 and Level-3 achieve significantly better performance on math benchmarks compared to augmentation methods. More importantly, as the number of generated questions per document increases to 150, our methods continue to show consistent improvements. In contrast, the performance of the rephrase and persona augmentation methods has already saturated as the dataset size expands. This demonstrates that compared with existing augmentation methods, our knowledge-guided methods more efficiently exploit limited reference documents by decomposing and recombining knowledge concepts, thereby enabling more scalable high-quality synthetic question generation. In previous experiments, we generated only 5 questions for Level-2 and 3 questions for Level-3. The results presented here indicate that our method has not yet reached its full potential in document utilization.
@@ 5.4 Ablation Study on Scaling Law Form
6 Conclusions and Future Work
In this work, we investigate the scaling laws of synthetic data and propose SynthLLM, a scalable web-scale data synthesis framework that systematically transforms pre-training data into high-quality synthetic datasets. Our findings show that synthetic data generated through SynthLLM consistently adheres to a rectified scaling law across various model sizes. By leveraging the fitted scaling law, we accurately predict model performance when the synthetic dataset size is doubled. These results suggest that systematically expanding synthetic datasets can continue to drive sustained and predictable performance improvements. Furthermore, our comparisons with other synthetic data generation and augmentation methods demonstrate that SynthLLM achieves superior performance and better scalability, providing a more effective approach for further enhancing model performance.
Our framework can be readily extended to other downstream domains, including code, physics and chemistry, and healthcare, expanding its applicability across diverse fields. Furthermore, SynthLLM is not limited to fine-tuning; we aim to explore its effectiveness in continued pre-training and even in the pre-training phase for synthetic data generation. Additionally, we will continue to refine and optimize our framework by developing more efficient strategies for leveraging pre-training data, with the goal of enhancing scaling efficiency of synthetic data.
Appendix A Used Prompts
@@
Here is an article crawl from the web, which our classifier has identified as having significant educational value for students learning math. Your task is to analyze this article and extract educational materials, specifically focusing on topics and key concepts that can enhance students’ understanding of mathematics and improve their problem-solving skills. Pay special attention to uncommon but important mathematical concepts that are crucial for a deeper understanding. ## Tasks 1. **Determine Educational Level:**       - Identify the appropriate educational level for the article based on its content. Choose from the       following options:         - Primary School         - Middle School         - High School         - College         - Graduate School         - Competition         - Other 2. **Identify Subject Area:**       - Specify the primary subject area of mathematics to which the article belongs (e.g., Calculus,       Geometry, Algebra, etc.). 3. **Extract Topics and Key Concepts:**       - **Topics:**         - List **1 to 5** main topics covered in the article.         - Use terms commonly recognized in academia or industry.       - **Key Concepts:**         - For each identified topic, list **5 to 20** related key concepts.         - Ensure these concepts are clearly articulated using standard academic or industry terms. ## Guidelines: - **Terminology:** Use precise and widely recognized academic or industry terminology for subjects, topics, and key concepts to maintain consistency and clarity. - **Educational Level Selection:** If appropriate, restrict the educational level to one of the following: "Primary School", "Middle School", "High School", "College", "Graduate School", or "Competition" to ensure accurate categorization. ## Text {{ text }} ## Output Format <level>Educational Level</level> <subject>Subject Area</subject> <topic> Topics: 1. topic 1 2. topic 2 </topic> <key_concept> Key Concepts: 1. topic 1:       1.1. key concept       1.2. key concept        2. topic 2:       2.1. key concept       … … </key_concept> ## Output
@@
Appendix B Examples about Extracted Topics and Key Concepts from Reference Document
In this section, we demonstrate some examples about extracted topics and key concepts from reference document by using previous prompt shown in Figure 11. We adopt a relative large Language model, Mistral-Large-Instruct-2407. The examples are shown in Figure 12 and 13. We can observe that Mistral-Large-Instruct-2407 is capable of extracting well-structured topics and their corresponding key concepts from the document.
<level>High School</level> <subject>Trigonometry</subject> <topic> Topics: 1. Trigonometric Functions and Identities 2. Geometry on a Sphere 3. Applications of Trigonometry 4. Complex Numbers and Trigonometry 5. Derivations and Proofs in Trigonometry </topic> <key_concept> Key Concepts: 1. Trigonometric Functions and Identities:       1.1. Sine, Cosine, and Tangent Functions       1.2. Trigonometric Identities (e.g., Pythagorean, Co-function, Sum and Difference)       1.3. Law of Sines and Law of Cosines       1.4. Inverse Trigonometric Functions       1.5. Trigonometric Equations and Their Solutions 2. Geometry on a Sphere:       2.1. Latitude and Longitude       2.2. Circumference of a Circle on a Sphere       2.3. Equation of a Circle on a Sphere       2.4. Spherical Trigonometry       2.5. Haversine Formula for Distance Calculation 3. Applications of Trigonometry:       3.1. Circular Motion and Angular Velocity       3.2. Calculating Angles in Right Triangles       3.3. Real-world Problems Involving Trigonometry (e.g., Crossing a Canyon)       3.4. Computing Sine and Cosine without a Calculator       3.5. Practical Uses of Trigonometric Functions in Engineering and Physics 4. Complex Numbers and Trigonometry:       4.1. Euler’s Formula and De Moivre’s Theorem       4.2. Complex Numbers and Trigonometric Functions       4.3. Roots of Unity and Their Relationship to Trigonometry       4.4. Inverse Hyperbolic Functions (e.g., Arcsinh)       4.5. Complex Numbers in Trigonometric Identities 5. Derivations and Proofs in Trigonometry:       5.1. Derivation of Trigonometric Identities       5.2. Proof of the Law of Sines and Law of Cosines       5.3. Derivation of Sum and Difference Formulas       5.4. Proof of the Haversine Formula       5.5. Derivation of Heron’s Formula </key_concept>

### v3

@@ 4.2 Main Experimental Results
SynthLLM Outperforms Other Synthetic Datasets. SynthLLM consistently surpasses other synthetic datasets across all benchmarks, demonstrating the effectiveness of SynthLLM. OpenMathInstruct-2 is a math dataset curated from MATH and GSM8K. Despite its larger size, the model trained on smaller dataset (SynthLLM-8B (3.2M)) achieves comparable performance on in-distribution test sets (MATH and GSM8K) while significantly outperforming it on OOD benchmarks, highlighting the superior generalization of SynthLLM. Additionally, our models consistently surpass Llama Instruct models of similar sizes and even match or outperform much larger models, such as NuminaMath-CoT-72B and Llama-3.1-70B-Instruct. Notably, we currently generate only five questions per document for Level-1 and Level-2 and three questions for Level-3. As shown in Figure 5, this is far from fully utilizing each document’s potential. Simply increasing the number of synthesized questions per document could further enhance benchmark performance.
@@ 4.3 Ablation Studies about Our SynthLLM Framework
Comparison with Direct Extraction-Based Synthesis. We compare Level-1 (direct extraction-based synthesis) with Level-2 and Level-3 across models of different sizes (1B and 3B). As shown in Table 4 of Appendix, on the 3B model, Level-2 and Level-3 significantly outperform Level-1 and the instruct-tuned version across all benchmarks. On the 1B model, Level-3 achieves substantial gains over Level-1 on most benchmarks, except for College and Gaokao. While each document yields five sampled questions, the final dataset size for Level-1 (2.3M) is notably smaller than that of Level-2 (2.6M). This is due to the high redundancy in direct extraction, where many duplicate questions are generated and later removed. This highlights the scalability limitations of direct extraction-based approach.
Diversity of Generated Questions. Compared to direct extraction-based synthesis, SynthLLM generates more diverse questions by decomposing and recombining knowledge concepts. To assess this, we compare Level-1 and Level-2 by measuring the similarity among generated questions within the same document. Specifically, we randomly select 2,000 documents and generate 50 questions per document for each method.
We then use all-MiniLM-L6-v2 (Wang et al., 2020) to obtain question embeddings and compute cosine similarity between each question pair from same $d$. The histogram of question similarity is shown in Figure 4. The results clearly indicate that Level-2 generates more diverse questions, as evidenced by the lower similarity scores among questions derived from same $d$.
Comparison with data augmentation methods on limited reference documents. We compare our Level-2 and Level-3 methods with augmentation techniques, including rephrase (Yu et al., 2024) and persona augmentation (Ge et al., 2024). We fix number of reference documents and evaluate different methods for scaling high-quality question generation from it. The effectiveness of each method will be measured by its ability to consistently improve model performance on downstream tasks as the dataset expands. Specifically, we randomly select 2,000 reference documents and generate 25, 50, to 150 questions per document using Level-2 and Level-3 methods, resulting in datasets ranging from 50K, 100K, to 300K questions. For rephrasing and persona augmentation, we need seed questions to conduct data augmentation. Therefore, we first adopt our Level-1 method to extract existing questions from the 2,000 fixed reference documents (yielding 6,500 questions) as the seed dataset. For above questions, we employ the same answer generator $\mathbf{M}^{A}$ (see Section 4.1). We train these datasets on the Llama-3.1-8B and present the results in Figure 5, where (a) shows MATH accuracy and (b) depicts average accuracy across benchmarks. Level-2 and Level-3 consistently outperform augmentation-based methods. Notably, as the number of generated questions per document increases to 150, our methods continue to yield performance gains, whereas rephrase and persona augmentation reach saturation. This highlights our approaches can efficiently utilize limited reference documents by decomposing and recombining knowledge concepts, enabling more scalable and high-quality question generation. In previous experiments, we generated only 5 questions per document for Level-2 and 3 questions per document for Level-3. These results suggest that our approach has not yet reached its full potential in document utilization and can be further expanded.
Figure 5: (a): The performances of Level-2, Level-3, and other augmentation methods on MATH benchmark. (b): The average performances across various benchmarks. The x axis denotes the sample number. The y axis represents the accuracy (%).
4.4 Preliminary Experiments in Coding Domain
To explore the generalizability and effectiveness of our method, we conduct preliminary experiments in the coding domain. We adopt the same SYNTHLLM framework (described in Section 3) and follow the experimental setup outlined in Section 4.1. Specifically, we generate 2.2M samples using our Level-1 baseline method, and 2.4M and 2.5M samples for our Level-2 and Level-3 methods, respectively. Mistral-Large-Instruct-2407 is used as the question generator, while Qwen2.5-Coder-32B-Instruct serves as the answer generator.
Model HumanEval MBPP
Llama-3.1-8B-Instruct $72.6$ $71.2$
Deepseek-Coder-v1.5-Instruct $75.6$ $73.4$
SynthLLM Level-1 $73.2$ $67.3$
SynthLLM Level-123 $78.7$ $72.5$
Table 3: HumanEval and MBPP Performance (%)
We then train the Llama-3.1-8B base model on these datasets. The results on HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) are summarized in Table 3. For comparison, we also include the performance of the instruction-tuned model (Llama-3.1-8B-Instruct) and a code-specific model (Deepseek-Coder-v1.5-Instruct (Guo et al., 2024)).
SynthLLM Level-1 represents the model trained solely on Level-1 data, while the Level-123 model denotes the model trained on the entire synthetic dataset (combined data from 3 methods). Our newly proposed Level-2 and Level-3 methods yield significant improvements in coding performance over the Level-1 baseline. SynthLLM Level-123 outperforms the instruction-tuned model and achieves performance comparable to the code-specific model, demonstrating the generalizability and effectiveness of our approach in coding.
5 Conclusions and Future Work
We investigate scaling laws of synthetic data by introducing SynthLLM. Our findings show that synthetic data generated by SynthLLM follows a rectified scaling law, enabling accurate performance prediction when doubling the dataset size. This suggests that systematically expanding the synthetic dataset yields sustained and predictable performance gains. Comparisons with existing synthetic data generation and augmentation methods demonstrate that SynthLLM offers superior performance and scalability. Our framework could extend to domains such as physics, law, and finance. Future directions include refining SynthLLM to optimize pre-training data utilization and answer verification for improved accuracy.
Appendix A Related Work
Synthetic Data Generation. Synthetic data has emerged as a promising method to mitigate data scarcity and reduce the costs associated with data collection and annotation. In this work, we focus specifically on synthetic data generation for the post-training phase. Conventional methods typically leverage a limited number of high-quality human-annotated seed examples to generate diverse augmentations via large language models (LLMs). Common techniques include: 1) sampling seed instructions as few-shot prompts for generating new instructions (Wang et al., 2022; Honovich et al., 2022; Toshniwal et al., 2025; Li et al., 2024a); and 2) rephrasing seed samples to create variations (Yu et al., 2024; Maini et al., 2024; Ge et al., 2024; Lambert et al., 2024; Luo et al., 2023). However, reliance on scarce high-quality seeds limits scalability and diversity. Recent approaches by Xu et al. (2024) and Li et al. (2024b) propose generating synthetic data from scratch by exploiting either LLM uncertainty or predefined knowledge taxonomies. In contrast, pre-training data—vast and inherently diverse—remains underutilized for scalable post-training synthetic data generation. Prior methods have either directly extracted samples from web documents (similar to our Level-1 approach) (Yue et al., 2024; Zhou et al., 2024; Yuan et al., 2025) or used document backtranslation (Li et al., 2024). Our proposed SynthLLM method aligns with this research direction but provides a more efficient framework for leveraging web documents to scale the diversity of synthetic data.
Scaling Law of LLMs. Scaling laws predict how model performance varies with model size and data volume, providing insights critical for optimizing computational resources during model training (Hestness et al., 2017; Rosenfeld et al., 2019; Kaplan et al., 2020; Hoffmann et al., 2022; Bahri et al., 2024). Recent advancements have introduced refined scaling laws, such as data-constrained scaling (Muennighoff et al., 2023), hyperparameter scaling Bi et al. (2024), and model distillation Busbridge et al. (2025). For fine-tuning LLMs, Hernandez et al. (2021) explored scaling behavior transitioning from unsupervised pre-training, emphasizing pre-training’s critical role. Lin et al. (2024) further introduced a rectified scaling law tailored specifically for fine-tuning tasks. Our work extends this rectified scaling law (Lin et al., 2024), demonstrating for the first time that scaling laws also apply robustly when fine-tuning LLMs using synthetic data.
Appendix B Details of Training Model and Our Dataset
Training Settings.
We employ Supervised Fine-Tuning (SFT) to train the model. We adopt AdamW optimizer (Loshchilov & Hutter, 2019) and set learning rate to $1\times 10^{-5}$ with a batch size of $512$. We utilize the linear learning rate scheduler and train the model for 3 epochs.
Our Dataset.
In this work, we focus on the math domain. After filtering high-quality documents from web data, we collect approximately $520\text{K}$ math-related documents. For Level-1 and Level-2, we generate 5 questions per document, followed by deduplication and decontamination (Yue et al., 2024; Toshniwal et al., 2025) using evaluation benchmarks, yielding around $2.3\text{M}$ and $2.6\text{M}$ samples, respectively. For Level-3, the constructed concept graph contains around 32 00032\,000 topics and 200 000200\,000 knowledge concepts. We conduct random walks on the topic sub-graph for five epochs, covering each topic in each epoch. For each sampled $\mathbf{K}^{g}$, we generate 3 questions, producing about $2.6\text{M}$ samples. Due to the high computational cost of training on large-scale datasets, we limit the number of sampled questions per document in Level-2 and constrain the number of random walk epochs in Level-3 to a relatively small scale (e.g., five epochs). It is still far from fully leveraging the potential of each document (see Figure 5). For $\mathbf{M}^{Q}$, the output temperature is $0.75$. For $\mathbf{M}^{Q}$, the output temperature of $\mathbf{M}^{A}$ is 0.
Statistics of Synthetic Questions and Answers.
Other Baselines.
We compare our synthetic data with other high-quality synthetic datasets. MAmmoTH2 (Yue et al., 2024), NaturalReasoning (Yuan et al., 2025), and JiuZhang3.0 (Zhou et al., 2024) employ a direct extraction-based synthesis approach on pre-training documents, similar to our Level-1 method. OpenMathInstruct-2 (Toshniwal et al., 2025) is a large-scale synthetic math dataset consisting of 14M samples, with GSM8K and MATH serving as the seed datasets. Additionally, we compare our dataset with NuminaMath (LI et al., 2024), a high-quality dataset that includes both synthetically generated data and extracted question-answer pairs from exam papers and mathematics discussion forums.
Furthermore, we evaluate our Level-2 and Level-3 approaches against existing methods designed to enhance question diversity through augmentation. Specifically, we include two popular augmentation approaches: the rephrase augmentation approach (Yu et al., 2024; Maini et al., 2024) and the persona augmentation approach (Ge et al., 2024; Lambert et al., 2024), both of which generate new questions based on the questions extracted using Level-1. For rephrase augmentation, we follow the same procedure as MetaMath (Yu et al., 2024), providing two-shot examples in the prompt to rephrase questions. For persona augmentation, we utilize the 200K personas released by the authors as prompts for question augmentation.
Evaluation.
To evaluate the math reasoning ability of our model trained on synthetic data, we utilize the evaluation code from Qwen-Math (Yang et al., 2024) to conduct assessments on several mainstream benchmarks: MATH (Hendrycks et al., 2021), GSM8K (Cobbe et al., 2021), OlympiadBench-Math (Olympiad) (He et al., 2024), CollegeMath (College) (Tang et al., ), Minerva Math (Minerva) (Lewkowycz et al., 2022), and Gaokao 2023 En (Gaokao) (Liao et al., 2024). The output temperature for all tested models is set to 0.
Computational Cost.
For generating questions, answers, and key concepts, we utilized a total of 32 Nvidia H100 GPUs, with each inference task running on 4 Nvidia H100 GPUs. The complete dataset generation process took approximately 3 days. For model training, we employed AMD MI300 GPUs; the largest training task utilized 32 AMD MI300 GPUs and required around 2 days to complete.
Appendix C More Results
In this section, we present additional experimental results about Section 4.
C.1 Ablation Study about Our Three Distinct Methods
We compare Level-1 (direct extraction-based synthesis) with Level-2 and Level-3 across models of different sizes (1B and 3B). As shown in Table 2 (Appendix), on the 3B model, Level-2 and Level-3 consistently outperform both Level-1 and the instruct-tuned version across all benchmarks. On the 1B model, Level-3 achieves substantial improvements over Level-1 on most benchmarks, except for College and Gaokao. Although each document contributes five sampled questions, the final dataset size for Level-1 (2.3M) is significantly smaller than that of Level-2 (2.6M). This discrepancy arises from the high redundancy in direct extraction, where many duplicate questions are generated and subsequently removed. These findings highlight the scalability limitations of direct extraction-based synthesis.
@@ C.1 Ablation Study about Our Three Distinct Methods
Table 4: Performance of 1B and 3B models on various math benchmarks. The number in $()$ represents the sample size. The best result is in bolded. All metrics are measured in percentage (%).
C.2 Ablation Study on Scaling Law Form
@@
C.3 Ablation Study on Answer Verification Process
The decision not to conduct further answer verification was motivated by our prior preliminary experiments, which showed that removing noisy samples during SFT yielded only marginal gains in model performance. We describe these experiments below. Our preliminary experiments followed the same experimental setup as the ablation studies described in Section 4.3. We randomly selected 2,000 reference documents and generated 50 questions per document using our Level-2 method. For each question, we generated 8 candidate answers.
Model MATH Average
Llama-3.1-8B wo filtering $42.0$ $36.5$
Llama-3.1-8B w filtering $42.2$ $36.6$
Table 5: MATH and Average Performance (%)
For each question, we used Llama-3.1-70B-Instruct as an answer evaluator to score its 8 candidate answers and discarded any judged "potentially incorrect". We then randomly picked one answer from the remaining candidates. Without filtering, one answer was randomly selected from all 8 candidates. In both cases, only one answer per question was retained for comparison. We then trained the Llama-3.1-8B base model on these generated datasets. We evaluate each trained model on the MATH benchmark and compute the average accuracy across all other benchmarks. As shown Table 5, filtering out noisy samples results in only marginal gains in model performance.
Recent studies (Tang et al., ; Li et al., 2025; Gandhi et al., 2025) report consistent findings. As shown in Table 7 of (Tang et al., ), adding a validation step for synthetic data does not improve performance on mathematical reasoning tasks. Table 2 of (Li et al., 2025) and Figure 6 of (Gandhi et al., 2025) further show that models trained even on entirely incorrect answers achieve performance comparable to those trained on correct answers. These results, along with our experiments, suggest that for SFT, model performance remains relatively robust to a certain degree of noise in answer correctness.
Appendix D Used Prompts
@@
Here is an article crawl from the web, which our classifier has identified as having significant educational value for students learning math. Your task is to analyze this article and extract educational materials, specifically focusing on topics and key concepts that can enhance students’ understanding of mathematics and improve their problem-solving skills. Pay special attention to uncommon but important mathematical concepts that are crucial for a deeper understanding. ## Tasks 1. **Determine Educational Level:**    - Identify the appropriate educational level for the article based on its content. Choose from the    following options:    - Primary School    - Middle School    - High School    - College    - Graduate School    - Competition    - Other 2. **Identify Subject Area:**    - Specify the primary subject area of mathematics to which the article belongs (e.g., Calculus,    Geometry, Algebra, etc.). 3. **Extract Topics and Key Concepts:**    - **Topics:**    - List **1 to 5** main topics covered in the article.    - Use terms commonly recognized in academia or industry.    - **Key Concepts:**    - For each identified topic, list **5 to 20** related key concepts.    - Ensure these concepts are clearly articulated using standard academic or industry terms. ## Guidelines: - **Terminology:** Use precise and widely recognized academic or industry terminology for subjects, topics, and key concepts to maintain consistency and clarity. - **Educational Level Selection:** If appropriate, restrict the educational level to one of the following: "Primary School", "Middle School", "High School", "College", "Graduate School", or "Competition" to ensure accurate categorization. ## Text {{ text }} ## Output Format <level>Educational Level</level> <subject>Subject Area</subject> <topic> Topics: 1. topic 1 2. topic 2 </topic> <key_concept> Key Concepts: 1. topic 1:    $1.$ key concept    1.2. key concept    … 2. topic 2:    2.1. key concept    … … </key_concept> ## Output
@@
Appendix E Examples about Extracted Topics and Key Concepts from Reference Document
In this section, we demonstrate some examples about extracted topics and key concepts from reference document by using previous prompt shown in Figure 10. We adopt a relative large Language model, Mistral-Large-Instruct-2407. The examples are shown in Figure 11 and 12. We can observe that Mistral-Large-Instruct-2407 is capable of extracting well-structured topics and their corresponding key concepts from the document.
<level>High School</level> <subject>Trigonometry</subject> <topic> Topics: 1. Trigonometric Functions and Identities 2. Geometry on a Sphere 3. Applications of Trigonometry 4. Complex Numbers and Trigonometry 5. Derivations and Proofs in Trigonometry </topic> <key_concept> Key Concepts: 1. Trigonometric Functions and Identities:    $1.$ Sine, Cosine, and Tangent Functions    1.2. Trigonometric Identities (e.g., Pythagorean, Co-function, Sum and Difference)    1.3. Law of Sines and Law of Cosines    1.4. Inverse Trigonometric Functions    1.5. Trigonometric Equations and Their Solutions 2. Geometry on a Sphere:    2.1. Latitude and Longitude    $2.$ Circumference of a Circle on a Sphere    2.3. Equation of a Circle on a Sphere    2.4. Spherical Trigonometry    2.5. Haversine Formula for Distance Calculation 3. Applications of Trigonometry:    3.1. Circular Motion and Angular Velocity    3.2. Calculating Angles in Right Triangles    $3.$ Real-world Problems Involving Trigonometry (e.g., Crossing a Canyon)    3.4. Computing Sine and Cosine without a Calculator    3.5. Practical Uses of Trigonometric Functions in Engineering and Physics 4. Complex Numbers and Trigonometry:    4.1. Euler’s Formula and De Moivre’s Theorem    4.2. Complex Numbers and Trigonometric Functions    4.3. Roots of Unity and Their Relationship to Trigonometry    $4.$ Inverse Hyperbolic Functions (e.g., Arcsinh)    4.5. Complex Numbers in Trigonometric Identities 5. Derivations and Proofs in Trigonometry:    5.1. Derivation of Trigonometric Identities    5.2. Proof of the Law of Sines and Law of Cosines    5.3. Derivation of Sum and Difference Formulas    5.4. Proof of the Haversine Formula    $5.$ Derivation of Heron’s Formula </key_concept>